# Improving Diffusion-based Data Augmentation with Inversion Spherical Interpolation

### v1

@@ 3.2 Inversion Interpolation
3.2.1 Inversion Pool
To get a faithful and diverse synthetic set by interpolating image pairs, we propose to conduct interpolation in the DDIM (Song et al., 2020) inversion space. There are two main motivations: 1) The sampling speed of DDIM is competitive due to the sampling of non-consecutive time steps. This can make our inverse process efficient. 2) We found that starting from the DDIM inversion can ensure a relatively high reconstruction result, especially conditioned on the learned concepts from Sec. 3.1.
The DDIM sampling has the following updating equation:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},c,t)}{\sqrt{\bar{\alpha}}_{t}})+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},c,t),$ (3)
where $x_{t}$ is the latent at timestep $t$ in reverse process. Based on Eq. (3), we can get the DDIM inversion update equation:
$x_{t}=\frac{\sqrt{\bar{\alpha}_{t}}}{(\sqrt{\bar{\alpha}_{t}}-1)}(x_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},c,t))+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},c,t)$ (4)
@@ 3.2.1 Inversion Pool
3.2.2 Random Spherical Interpolation
Figure 4: Spherical Interpolation
Since Gaussian noises are received as input during the training process of the diffusion model, we need to ensure the initial noise for the denoising process also resides in a Gaussian distribution. Since each inversion in $\mathcal{I}^{i}$ is in a Gaussian distribution, we propose to conduct spherical interpolation (Shoemake, 1985) on the inversion pairs. This can maintain the interpolation result in Gaussian distribution, and it then can be the initial noise for the denoising process.
After getting the inversion set $\mathcal{I}^{i}$, we randomly select two DDIM inversions $I_{a},I_{b}$ (ignored the superscript) as a pair from $\mathcal{I}^{i}$. For this pair, we conduct the random spherical interpolation. As shown in Figure 4, our random spherical interpolation consists of interpolation (Green arrow) and extrapolation (Blue arrow). The images are the final denoised results. Below are the details of interpolation and extrapolation:
Interpolation. The interpolation can be expressed as follows:
$Z=\frac{sin((1-\lambda)\alpha)}{sin(\alpha)}I_{a}+\frac{sin(\lambda\alpha)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,1],$ (5)
where $\lambda$ is the interpolation strength, $\alpha=arccos(\frac{I_{a}^{T}I_{b}}{(||I_{a}||||I_{b}||))}$ and $Z$ is the final interpolation result.
Extrapolation. In order to increase the interpolation range, we also leverage extrapolation:
$Z=\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}I_{a}-\frac{sin(\lambda\alpha)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,2\pi/\alpha-1]$ (6)
Unified Interpolation. According to the periodicity of trigonometric functions, we can get a unified representation of interpolation and extrapolation by combining Eq. (5) and Eq. (6) as follows:
$Z=\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}I_{a}-\frac{sin(\lambda\alpha)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,2\pi/\alpha],$ (7)
where $\lambda$ is a random value range in a full interpolation cycle. It can decide the interpolation type (interpolation or extrapolation) and control the relative distance between $Z$ and $I_{a}$, $I_{b}$. According to the expansion rate of the $i_{th}$ category, we repeat the sampling and interpolation. Then, we collect all the interpolation results into $\mathcal{Z}^{i}$, which will be used as the initial noises in Sec. 3.3.
@@ 3.3 Two-stage Denoising
In order to further increase the diversity of synthetic images, we design a two-stage denoising process (c.f., Figure 5). We split the denoising process into two stages with a split ratio $s\in[0,1]$. The first stage includes time steps from $T$ to $sT$. The second stage includes time steps from $sT$ to 00. The main difference between the two stages is the prompt used.
Suffixed Prompt. For a specific dataset, we will generate a few suffixes that can summarize the context of this dataset. First, we input each $X_{j}\in\mathcal{O}$ into a pre-trained vision language model (VLM) (e.g., BLIP (Li et al., 2022)) to get the corresponding caption. After getting all the captions, we input them into a large language model (LLM) (e.g., GPT-4 (Achiam et al., 2023)) to summarize them into a few descriptions with the following format: “a photo of a [metaclass] [suffix]”. Thus, we can get a few suffixes for a dataset2. Based on the captioning ability of VLM and the powerful generalization ability of LLM, these suffixes summarize the high-frequency context in the dataset. For each $Z\in\mathcal{Z}^{i}$, we randomly sample one suffix then concat the plain prompt with this suffix into: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass] [suffix]”.
@@
$x_{t-1}=\sqrt{\bar{\alpha_{t-1}}}(\frac{x_{t}-\sqrt{1-\bar{\alpha_{t}}}\epsilon_{\theta}}{\sqrt{\bar{\alpha_{t}}}})+\sqrt{1-\bar{\alpha_{t-1}}}\epsilon_{\theta}\quad\left\{\begin{array}[]{lr}\epsilon_{\theta}=\epsilon_{\theta}(x_{t},c^{*},t),&t\in(sT,T]\\ \epsilon_{\theta}=\epsilon_{\theta}(x_{t},c,t),&t\in[0,sT)\end{array}\right.,$ (8)
@@ 4.1 Few-shot classification
Real-Filter (He et al., 2022) 55.03 30.40 87.66 57.50 57.65$(+3.09)$ 67.36 47.21 90.14 69.64 70.66$(-0.45)$
Real-Guidance (He et al., 2022) 54.67 36.27 86.22 42.45 54.90$(+0.34)$ 69.81 54.77 90.06 70.97 71.40$(+0.29)$
Real-Mix (Wang et al., 2024) 51.00 22.53 88.63 52.06 53.56$(-1.00)$ 64.65 44.16 89.40 75.72 68.48$(-2.63)$
Da-Fusion (Trabucco et al., 2023) 59.40 34.98 88.64 51.90 58.73$(+4.17)$ 72.05 51.14 90.47 77.61 72.82$(+1.71)$
Diff-AUG (Wang et al., 2024) 61.14 39.48 89.24 62.28 63.04$(+8.78)$ 72.02 55.43 90.28 81.65 74.85$(+3.74)$
Diff-Mix (Wang et al., 2024) 56.18 32.61 88.77 56.39 58.49$(+3.93)$ 70.16 52.33 90.79 78.68 72.99$(+1.88)$
@@
Real-Filter (He et al., 2022) 73.66 31.44 90.85 72.51 67.12$(+3.05)$ 82.19 46.43 93.14 84.69 76.61$(-0.65)$
Real-Guidance (He et al., 2022) 74.74 35.10 91.23 62.34 65.85$(+1.78)$ 83.24 53.15 93.22 81.97 77.90$(+0.64)$
Real-Mix (Wang et al., 2024) 72.25 32.91 90.92 70.55 66.66$(+2.59)$ 80.95 47.01 93.16 83.78 76.23$(-1.03)$
Da-Fusion (Trabucco et al., 2023) 76.24 34.20 93.03 71.07 68.64$(+4.57)$ 83.97 51.68 93.68 85.00 78.58$(+1.32)$
Diff-AUG (Wang et al., 2024) 77.24 40.08 92.27 76.48 71.52$(+7.45)$ 84.10 56.24 93.93 87.59 80.47$(+3.21)$
Diff-Mix (Wang et al., 2024) 74.49 36.48 92.29 73.01 69.07$(+5.00)$ 81.98 53.92 93.98 85.42 78.83$(+1.57)$
@@
Table 1: Few-shot classification. 5-shot and 10-shot results on four fine-grained datasets with two backbones. “Original” means the model trained on the original set without DA. Green and red numbers are increase and decrease values w.r.t. “Original”. All results are averaged on three trials.
Settings. To evaluate the Diff-II’s augmentation capacity based on few samples, we conducted few-shot classification on four domain-specific fine-grained datasets: Caltech-UCSD Birds (Wah et al., 2011), FGVC-Aircraft (Maji et al., 2013), Stanford Cars (Krause et al., 2013) and Oxford Pet (Parkhi et al., 2012), with shot numbers of 5, 10. We used the augmented datasets to fine-tune two backbones: 224$\times$224-resolution ResNet50 (He et al., 2016) pre-trained on ImageNet1K (Deng et al., 2009) and 384$\times$384 ViT-B/16 (Dosovitskiy et al., 2020) pre-trained on ImageNet21K. We compared our method with six diffusion-based augmentation methods: Real-Filter, Real-Guidance (He et al., 2022), Da-Fusion (Trabucco et al., 2023), Real-Mix, Diff-AUG and Diff-Mix (Wang et al., 2024). We fix $s$ to 0.3 for 5-shot and 0.1 for 10-shot. For fairness, the expansion rate was 5 for all methods. For the classifier training process, we followed the joint training strategy of  (Trabucco et al., 2023): replacing the data from the original set with synthetic data in a replacement probability during training. We fixed the replacement probability with 0.5 for all methods. More details are in the Appendix A.3.
Results. From the results in Table 1, we have several observations: 1) Compared with training on the original set, our method can improve the average accuracy from $3.56\%$ to $10.05\%$, indicating that our methods can effectively augment domain-specific fine-grained datasets. 2) Our method can outperform all the comparison methods in all settings, demonstrating the effectiveness of our method for few-shot scenarios. 3) Our method achieves greater gains for smaller shots (i.e., 5shot) and weaker backbone (i.e., ResNet50), showing our method is robust to challenging settings.
@@ 4.2 Long-tail classification
Settings. To evaluate the Diff-II’s augmentation capacity for datasets with imbalanced samples, we experimented with our methods on the long-tail classification task. Following the previous settings (Cao et al., 2019; Liu et al., 2019; Park et al., 2022; Wang et al., 2024), we evaluated our method on two domain-specific long-tail datasets: CUB-LT (Samuel et al., 2021) and Flower-LT (Wang et al., 2024), with imbalance factor (IF) of 100, 20,10. We used the 224$\times$224-resolution ResNet50 (mentioned in Sec. 4.1) as the backbone. We fixed $s$ to 1.0 for all settings. We compared our method with five methods: oversampling-based CMO (Park et al., 2022), re-weighting CMO (CMO+DRW (Cao et al., 2019)), diffusion-based Real-Filter, Real-Guidance, Real-Mix, and Diff-Mix. For fairness, the expansion rate was 5, and the replacement probability was 0.5 for all diffusion-based methods. More details are in the Appendix A.3.
@@
CMO (Park et al., 2022) 32.94 44.08 57.62 44.88$(-0.65)$ 83.95 91.43 95.19 90.19$(+1.40)$
CMO-DRW (Cao et al., 2019) 32.57 46.43 59.25 46.08$(+0.55)$ 84.07 92.06 95.92 90.68$(+1.89)$
Real-Gen (Wang et al., 2024) 45.86 53.43 61.42 53.57$(+8.04)$ 83.56 91.84 95.22 90.21$(+1.42)$
Real-Mix (Wang et al., 2024) 47.75 55.67 62.27 55.23$(+9.70)$ 85.19 92.96 96.04 91.40$(+2.61)$
Diff-Mix (Wang et al., 2024) 50.35 58.19 64.48 57.67$(+12.14)$ 89.46 93.86 96.63 93.32$(+4.53)$
Ours 51.21 62.31 70.28 61.27$(+15.74)$ 89.54 94.39 97.35 93.76$(+4.97)$
Table 2: Long-tail classification results on CUB-LT and Flower-LT. “CE” is a plain baseline that trains a classifier on the original set with the Cross-Entropy loss. It contains no operations designed for long-tail tasks. “IF” is the imbalanced factor, where a larger IF indicates more imbalanced data distribution. Green and red numbers are the increase and decrease values w.r.t. CE. “Ours” results are averaged on three trials, and other results are from (Wang et al., 2024)
@@
4.3 Out-of-Distribution classification
Settings. To evaluate whether the synthetic data generated by Diff-II can benefit the generalization capacity of the classifier, we conducted out-of-distribution (OOD) classification experiments. To be specific, we trained a 224$\times$224-resolution ResNet50 (c.f., Sec. 4.1) with the original set of 5shot CUB and corresponding synthetic data (same with Sec. 4.1). Then we tested it on an out-of-distribution dataset: Waterbird (Sagawa et al., 2019). The Waterbird is constructed in this way: segment the CUB’s foregrounds and paste them into the images from Places (Zhou et al., 2017). The images from Places provide new backgrounds for CUB’s foregrounds. The Waterbird dataset can be divided into 4 groups: (land, land), (water, water), (land, water) and (water, land). The former in brackets refers to the type of foreground bird (water bird or land bird), and the latter refers to the type of background bird (water background or land background). For example, the (land, water) indicates the land bird in the water background. Besides, the comparison methods were six diffusion-based data augmentation methods: Real-Filter, Real-Guidance, Da-Fusion, Real-Mix, Diff-AUG and Diff-Mix. We used the same hyper-parameters with Sec. 4.1.
Results. As shown in Table 3, we can have two observations: 1) Our method can significantly improve the classification ability of the classifier on the background-shift out-of-distribution dataset by augmenting the original dataset. For example, the average accuracy can be improved by $11.39\%$ compared to the “Original” (no augmentation) one. This shows that the data generated by our Diff-II has good diversity, so it is possible to train a classifier that is robust to the background. 2) Our method can outperform all the comparison methods in 4 groups. Especially in (water, land) group, Diff-II can outperform the second-best method (Diff-AUG) by $3.45\%$. This demonstrates the excellent ability of our method to generate faithful and diverse images.

### v2

@@ 3.2 Inversion Interpolation
3.2.1 Inversion Pool Construction
To get a faithful and diverse synthetic set by interpolating image pairs, we propose to conduct interpolation in the DDIM [32] inversion space [^5] More backgrounds about DDIM and DDIM inversion are in appendix. There are two main motivations: 1) The sampling speed of DDIM is competitive due to the sampling of non-consecutive time steps. This can make our inverse process efficient. 2) We found that starting from the DDIM inversion can ensure a relatively high reconstruction result, especially conditioned on the learned concepts from Sec. 3.1.
DDIM sampling has the following updating equation:
$\begin{split}x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},c,t)}{\sqrt{\bar{\alpha}}_{t}})\\ +\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t},c,t),\end{split}$ (3)
where $x_{t}$ is the latent at timestep $t$ in reverse process. Based on Eq. (3) and $\theta(x_{t},c,t)\simeq\theta(x_{t-1},c,t)$ , we can get the DDIM inversion update equation:
$\begin{split}x_{t}\simeq\frac{\sqrt{\bar{\alpha}_{t}}}{(\sqrt{\bar{\alpha}_{t}}-1)}(x_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}(x_{t-1},c,t))\\ +\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t-1},c,t)\end{split}$ (4)
@@ 3.2.1 Inversion Pool Construction
3.2.2 Random Circle Interpolation
Since Gaussian noises are received as input during the training process of the diffusion model, we need to ensure the initial noise for the denoising process also resides in a Gaussian distribution. Since each inversion in $\mathcal{I}^{i}$ is in a Gaussian distribution, the common linear interpolation will lead to a result that is not in Gaussian distribution. Thus, we propose to conduct circle interpolation on the inversion pairs. This operation has a larger interpolation range (which increases the diversity) and can maintain the interpolation result in Gaussian distribution3. Thus, it can be the initial noise for the denoising process.
After getting the inversion set $\mathcal{I}^{i}$, we randomly select two DDIM inversions $I_{a},I_{b}$ (ignored the superscript) as a pair from $\mathcal{I}^{i}$. For this pair, we conduct the random circle interpolation.
Circle interpolation. The circle interpolation can be intuitively understood as rotating from one to another and it can be expressed as follows:
$Z=\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}I_{a}-\frac{sin(\lambda\alpha)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,\frac{2\pi}{\alpha}],$ (5)
where $\alpha=arccos(\frac{I_{a}^{T}I_{b}}{(||I_{a}||||I_{b}||))})$ and $Z$ is the interpolation result. $\lambda$ is a random interpolation strength, which can decide the interpolation type (interpolation or extrapolation) and control the relative distance between $Z$ and $I_{a}$, $I_{b}$.
As shown in Figure 4, the path of circle interpolation is the circle composed of the Green Arc and the Blue Arc. According to the rotation direction, we can decompose the circle interpolation into spherical interpolation and spherical extrapolation [31]:
Figure 4: An illustration for the proposed circle interpolation.
Spherical Interpolation. The spherical interpolation means rotate along the shortest path (cf., the Green Arc of Figure 4) and it can be expressed as follows:
$Z=\frac{sin((1-\lambda)\alpha)}{sin(\alpha)}I_{a}+\frac{sin(\lambda\alpha)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,1]$ (6)
Spherical Extrapolation. The spherical extrapolation means rotate along the opposite direction of the interpolation path (cf., the Blue Arc of Figure 4) and it can be expressed as follows:
$Z=\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}I_{a}-\frac{sin(\lambda\alpha)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,\frac{2\pi}{\alpha}-1]$ (7)
According to the periodicity of trigonometric functions, we can see that Eq. (5) is a unified representation of spherical interpolation (Eq. (6)) and spherical extrapolation (Eq. (7)). Based on the expansion rate of the $i_{th}$ category, we repeat the sampling and interpolation. Then, we collect all the interpolation results into $\mathcal{Z}^{i}$, which will be used as the initial noises in Sec. 3.3.
@@ 3.3 Two-stage Denoising
In order to further increase the diversity of synthetic images, we design a two-stage denoising process (cf., Figure 5). We split the denoising process into two stages with a split ratio $s\in[0,1]$. The first stage includes time steps from $T$ to $sT$. The second stage includes time steps from $sT$ to 00. The main difference between the two stages is the prompt used.
Suffixed Prompt. For a specific dataset, we will generate a few suffixes that can summarize the context of this dataset. First, we input each $X_{j}\in\mathcal{O}$ into a pre-trained vision language model (VLM) (e.g., BLIP [18]) to get the corresponding caption. After getting all the captions, we input them into a large language model (LLM) (e.g., GPT-4 [1]) to summarize them into a few descriptions with the following format: “a photo of a [metaclass] [suffix]”. Thus, we can get a few suffixes for a dataset3. Based on the captioning ability of VLM and the powerful generalization ability of LLM, these suffixes summarize the high-frequency context in the dataset. For each $Z\in\mathcal{Z}^{i}$, we randomly sample one suffix then concat the plain prompt with this suffix into: “a photo of a $[V_{1}^{i}]$ $[V_{2}^{i}]$$[V_{n}^{i}]$ [metaclass] [suffix]”.
@@
$\begin{split}x_{t-1}=\sqrt{\bar{\alpha_{t-1}}}(\frac{x_{t}-\sqrt{1-\bar{\alpha_{t}}}\epsilon_{\theta}}{\sqrt{\bar{\alpha_{t}}}})+\sqrt{1-\bar{\alpha_{t-1}}}\epsilon_{\theta}\quad\end{split}$ (8)
@@ 4.1 Few-shot Classification
Mixup[39] 54.33 36.75 86.36 41.43 54.72$(+0.16)$ 70.00 55.28 89.49 70.28 71.26$(+0.15)$
CutMix[38] 51.74 33.09 86.52 36.19 51.89$(-2.67)$ 68.91 51.05 89.71 67.71 69.35$(-1.76)$
Real-Filter [12] 55.03 30.40 87.66 57.50 57.65$(+3.09)$ 67.36 47.21 90.14 69.64 70.66$(-0.45)$
Real-Guidance [12] 54.67 36.27 86.22 42.45 54.90$(+0.34)$ 69.81 54.77 90.06 70.97 71.40$(+0.29)$
Real-Mix [36] 51.00 22.53 88.63 52.06 53.56$(-1.00)$ 64.65 44.16 89.40 75.72 68.48$(-2.63)$
Da-Fusion [33] 59.40 34.98 88.64 51.90 58.73$(+4.17)$ 72.05 51.14 90.47 77.61 72.82$(+1.71)$
Diff-AUG [36] 61.14 39.48 89.24 62.28 63.04$(+8.78)$ 72.02 55.43 90.28 81.65 74.85$(+3.74)$
Diff-Mix [36] 56.18 32.61 88.77 56.39 58.49$(+3.93)$ 70.16 52.33 90.79 78.68 72.99$(+1.88)$
@@
Mixup[39] 76.07 36.15 90.77 61.92 66.23$(+2.16)$ 85.24 52.94 92.81 82.46 78.36$(+1.10)$
CutMix[38] 72.97 32.79 90.22 59.16 63.79$(-0.28)$ 83.95 51.56 92.59 81.71 77.45$(+0.19)$
Real-Filter [12] 73.66 31.44 90.85 72.51 67.12$(+3.05)$ 82.19 46.43 93.14 84.69 76.61$(-0.65)$
Real-Guidance [12] 74.74 35.10 91.23 62.34 65.85$(+1.78)$ 83.24 53.15 93.22 81.97 77.90$(+0.64)$
Real-Mix [36] 72.25 32.91 90.92 70.55 66.66$(+2.59)$ 80.95 47.01 93.16 83.78 76.23$(-1.03)$
Da-Fusion [33] 76.24 34.20 93.03 71.07 68.64$(+4.57)$ 83.97 51.68 93.68 85.00 78.58$(+1.32)$
Diff-AUG [36] 77.24 40.08 92.27 76.48 71.52$(+7.45)$ 84.10 56.24 93.93 87.59 80.47$(+3.21)$
Diff-Mix [36] 74.49 36.48 92.29 73.01 69.07$(+5.00)$ 81.98 53.92 93.98 85.42 78.83$(+1.57)$
@@
Table 1: Few-shot classification. 5-shot and 10-shot results (averaged on three trials) on four fine-grained datasets with two backbones. “Original” means the model trained on the original set w/o DA. Green and red numbers are increase and decrease values w.r.t. “Original”.
Settings. To evaluate the Diff-II’s augmentation capacity based on few samples, we conducted few-shot classification on four domain-specific fine-grained datasets: CUB [35], Aircraft [20], Cars [16] and Pet [27], with shot numbers of 5, 10. We used the augmented datasets to fine-tune two backbones: 224$\times$224-resolution ResNet-50 [11] pre-trained on ImageNet1K [5] and 384$\times$384 ViT-B/16 [7] pre-trained on ImageNet21K. We compared our method with two DA augmentation methods: Mixup [39], CutMix [38] and six diffusion-based DA methods: Real-Filter, Real-Guidance [12], Da-Fusion [33], Real-Mix, Diff-AUG and Diff-Mix [36]. We fixed $s$ to 0.3 for 5-shot and 0.1 for 10-shot. For fairness, the expansion rate was 5 for all methods. For the classifier training process, we followed the joint training strategy of  [33]: replacing the data from the original set with synthetic data in a replacement probability during training. We fixed the replacement probability with 0.5 for all methods. More details are in the Appendix.
Results. From the results in Table 1, we have several observations: 1) Compared with training on the original set, our method can improve the average accuracy from $3.56\%$ to $10.05\%$, indicating that our methods can effectively augment domain-specific fine-grained datasets. 2) Our method can outperform all the comparison methods in all settings, demonstrating the effectiveness of our method for few-shot scenarios. 3) Our method achieves greater gains for smaller shots (i.e., 5-shot) and weaker backbone (i.e., ResNet-50), showing our method is robust to challenging settings.
@@ 4.2 Long-tail Classification
Settings. To evaluate the Diff-II’s augmentation capacity for datasets with imbalanced samples, we experimented with our methods on the long-tail classification task. Following the previous settings [3, 19, 26, 36], we evaluated our method on two domain-specific long-tail datasets: CUB-LT [30] and Flower-LT [36], with imbalance factor (IF) of 100, 20, and 10. We used the 224$\times$224-resolution ResNet-50 (mentioned in Sec. 4.1) as the backbone. We fixed $s$ to 1.0 for all settings. For those categories with only one image, we randomly sample a noise for subsequent generation since we can not interpolate. We compared our method with five methods: oversampling-based CMO [26], re-weighting CMO [3], diffusion-based Real-Filter, Real-Guidance [12], Real-Mix, and Diff-Mix [36]. For fairness, the expansion rate was 5, and the replacement probability was 0.5 for all diffusion-based methods. More details are in the Appendix.
@@
CMO [26] 32.94 44.08 57.62 $-0.65$ 83.95 91.43 95.19 $+1.40$
CMO-DRW [3] 32.57 46.43 59.25 $+0.55$ 84.07 92.06 95.92 $+1.89$
Real-Gen [36] 45.86 53.43 61.42 $+8.04$ 83.56 91.84 95.22 $+1.42$
Real-Mix [36] 47.75 55.67 62.27 $+9.70$ 85.19 92.96 96.04 $+2.61$
Diff-Mix [36] 50.35 58.19 64.48 $+12.14$ 89.46 93.86 96.63 $+4.53$
Ours 51.21 62.31 70.28 $+15.74$ 89.54 94.39 97.35 $+4.97$
Table 2: Long-tail classification results on CUB-LT and Flower-LT. “CE” is a plain baseline that trains a classifier on the original set with the Cross-Entropy loss. It contains no operations designed for long-tail tasks. “IF” is the imbalanced factor, where a larger IF indicates more imbalanced data distribution. Green and red numbers are the increase and decrease values w.r.t. CE. “Ours” results are averaged on three trials, and other results are from [36]
@@
4.3 Out-of-Distribution (OOD) Classification
Settings. To evaluate whether the synthetic data generated by Diff-II can benefit the generalization capacity of the classifier, we conducted OOD classification experiments. To be specific, we trained a 224$\times$224-resolution ResNet-50 (cf., Sec. 4.1) with the original set of 5-shot CUB and corresponding synthetic data (same with Sec. 4.1, and evaluated it on a prevalent OOD dataset: Waterbird [28]. Besides, the comparison methods were six diffusion-based data augmentation methods: Real-Filter [12], Real-Guidance [12], Da-Fusion [33], Real-Mix [36], Diff-AUG [36] and Diff-Mix [36]. We used the same hyper-parameters with Sec. 4.1.
Results. As shown in Table 4.4, we can have two observations: 1) Our method can significantly improve the classification ability of the classifier on the background-shift out-of-distribution dataset by augmenting the original dataset. For example, the average accuracy can be improved by $11.39\%$ compared to the “Original” (no augmentation) one. This shows that the data generated by our Diff-II has good diversity, so it is possible to train a classifier that is robust to the background. 2) Our method can outperform all the comparison methods in 4 groups. Especially in (water, land) group, Diff-II can outperform the second-best method (Diff-AUG) by $3.45\%$. This demonstrates the excellent ability of our method to generate faithful and diverse images.